OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -7, an = 4 • an-1

hero (hero):

\(a_1 = -7\) \(a_n = 4(a_{n - 1})\) Is that what you have written @Loveiskey18

hero (hero):

If so, then \(a_2 = 4(a_{2 -1})\) \(=4(a_1)\) \(= 4(-7) \) \(= -28\)

OpenStudy (anonymous):

Yes, my apologies I was doing something else. Afterwards... -28, -112, -448, -1792, -7168, -28,672

hero (hero):

\(-7\) is the first term. You don't need \(-28,672\)

hero (hero):

I'm also assuming your work is correct.

OpenStudy (anonymous):

-7, -28, -112, -448, -1792, -7168

OpenStudy (anonymous):

Okay, Thank You.